If 4 less than 3 times a certain number is 2 more than the number, what is the number?
Answer by AAfter.Com(18)

Let the given number be n.
Then, 3n - 4 = n + 2
=> 2n = 6
=> n = 3
Hence, the given number is 3.
Check: 3 X 3 - 4 = 5
3 + 2 = 5